OpenStudy (anonymous):

How do you use y=a(x-h)^2+k Please help I'll give medals

OpenStudy (campbell_st):

this is the vertex form of a parabola... if you have a parabola \[y = x^2 + 6x + 5\] complete the square in x \[y = (x^2 + 6x + 9) + 5 - 9\] so factoring \[y = (x + 3)^2 - 4\] so the parabola has a vertex at (-3, 4) hope it makes sense